OpenStudy (anonymous):

tan theta x cos theta = 3 verify the identity

OpenStudy (anonymous):

it is not an identity

OpenStudy (anonymous):

in fact, \[\tan(\theta)\cos(\theta)=\sin(\theta)\] and sine is never 3

OpenStudy (anonymous):

I thought I knew what I was doing until I got to this identity. Thank you.